$SPAD/src/input d01apf.input

The Axiom Team December 3, 2016

Abstract

Contents

2

1 License

— license —

--Copyright The Numerical Algorithms Group Limited 1994.

———-

— * —

\getchunk{license}
)clear all
showArrayValues true
showScalarValues true
e:Expression Float:=sin(10*X)
f:ASP1(G):=retract e
a:SF:=1.0e-6
b:SF:=1.0
alpha:SF:=-0.5
beta:SF:=-0.5
epsabs:SF:=0.0
epsrel:SF:=1.0e-4
key:=1
liw:=200
lw:=4*liw
result:=d01apf(a,b,alpha,beta,key,epsabs,epsrel,lw,liw,-1,f)

———-

References^{[1] nothing}

4